OpenStudy (anonymous):

Estimate the radius of the circle with given circumference round answer to the nearest 10th 80km

OpenStudy (anonymous):

C=2*pi*r r=C/(2*pi) r=80km/(2*pi)

OpenStudy (campbell_st):

well (80/6) is a good estimate... using 3 as an estimate for pi

OpenStudy (anonymous):

so confused so the answer is 480

OpenStudy (anonymous):

when they say estimate the circumference, do you tell you use to use 3.14 as pi?

OpenStudy (campbell_st):

the question is estimate the RADIUS of a circle with circumference of 80 units... the diameter is 80/3 which is 26 2/3 units... halve this to get the radius (26 2/3) = 13 1/3 so the radius is approx 13.3 units